OpenStudy (anonymous):

Which of the following sample will have the greatest average speed at 355 K? CH4 Ne C2H4 Cl2

OpenStudy (chmvijay):

speed inversely proportional to the molecular mass higher molecular weight lower is speed

OpenStudy (chmvijay):

methane

OpenStudy (anonymous):

average speed=\[\sqrt{\frac{ 8RT }{ \pi M}}\] M=Molecular Mass is lowest for CH4 Hence, it has highest average speed.